### v1

# Multi-height analysis of rational points of toric varieties

Abstract.
We study the multi-height distribution of rational points of smooth, projective and split toric varieties over $\mathbf{Q}$ using the lift of the number of points to universal torsors.
2020 Mathematics Subject Classification:
Primary NT11D45, AG14M25
For a quasi-Fano variety $V$ with infinitely many rational points, one may asymptotically study the finite set of rational points with a bounded height $H$ associated with the anticanonical line bundle $\omega_{V}^{-1}$. In [11], [3] and [13], Batyrev, Franke, Manin, Tschinkel and Peyre gave strong evidence supporting conjectures relating the asymptotic behavior of the number of points of bounded height on open subsets of $V$ to geometrical invariants of $V$.
In the case of toric varieties, the asymptotic behavior was studied in [1] and [2] using harmonic analysis on adelic groups. In [18], Salberger was the first to explicitly use universal torsors in relation to points of bounded height. In particular, he was able to give a new proof of the theorem of Batyrev and Tschinkel for smooth projective split toric varieties over $\mathbf{Q}$. His work highlighted how the constant in the conjectural asymptotic formula is related to the product of local densities over universal torsors. This was later studied in a more general setting by Peyre in [14] and [15].
More recently in [17], Peyre proposed a new framework to study the asymptotic distribution of rational points of bounded height. Instead of considering a height relative to one line bundle, it is natural to consider all possible heights. The purpose of this paper is to study the multi-height distribution of rational points of smooth, projective and split toric varieties over $\mathbf{Q}$. We prove that the expected asymptotic behavior holds for these varieties. In particular, we show that there is no accumulating subset for the multi-height distribution in the toric case. This is because we consider the rational points whose multi-height are in the interior of dual of the effective cone.
A result concerning this all heights approach was given in [10], using height zeta functions. It seems that the condition $\beta=\pi^{\vee}(u)$ where $u\in(C_{\operatorname{eff}}(X)^{\vee})^{\circ}$ is missing which would make it possible to construct counterexamples to their main result. Our method using descent is totally different. It is inspired by the work of Salberger in [18] and uses the method of counting the number of points of a lattice in a bounded domain developed by [8].
The author thanks his PhD advisor Emmanuel Peyre for all the remarks and suggestions he made during the writing process of this article.
1. The Manin-Peyre conjecture for the multi-height distribution of rational points on Fano varieties
1.1. Preliminaries
In this paragraph, we recall, following Peyre [17], how to define a system of Arakelov heights for a smooth, projective and geometrically integral variety $V$ over a number field $K$.
Notations.
We denote by $M_{K}$ the set of places of $K$, by $M_{K}^{0}$ the set of finite places and by $M_{K}^{\infty}$ its infinite places.
If $v\in M_{K}^{0}$, we write $q_{v}=\sharp\mathbf{F}_{v}$ the cardinal of the residual field associated to $v$, while if $v\in M_{K}^{\infty}$, we set $q_{v}=\exp(-[K_{v}:\mathbf{R}])$.
To $w\in M_{K}$ which divides $v\in M_{\mathbf{Q}}$, we define $|-|_{w}$ over $K_{w}$ by setting for any $x\in K_{w}$ :
$|x|_{w}=|N_{K_{w}/\mathbf{Q}_{v}}(x)|_{v}.$
It is an absolute value except when $K_{w}$ is $\mathbf{C}$ in which case it is the square of an absolute value.
To an inversible sheaf $\mathcal{L}$ over a scheme $T$, we associate the line bundle $L$ whose sections correspond to $\mathcal{L}^{\vee}$. We denote $L^{\times}=L\setminus z(T)$, where $z(T)$ is the closed subscheme given by the zero section.
We begin by defining the notion of an adelic norm over a vector bundle.
Definition 1.1.
[…]
For any field extension $L$ over $K$, for any $P\in V(L)$, we denote by $E_{P}\subset E(L)$ the $L$-vector space which corresponds to the fiber $e^{-1}(P)$ of $e$ over $P$.
An adelic norm $E$ over $e$ is the data of a family of continuous applications $(||-||)_{v\in M_{K}}$
$||-||_{v}:E(K_{v})\rightarrow\mathbf{R}_{+}$
such that:
  1. (1)
    If $v\in M_{K}^{0}$, for any $P\in V(K_{v})$, the restriction of $||-||_{v}$ to $E_{P}$ defines an ultrametric norm with values in the image of the absolute value $|-|_{v}$ i.e $q_{v}^{\mathbf{Z}}$.
  2. (2)
    If $v$ is a real place, for any $P\in V(K_{v})$, the restriction to $E_{P}$ is an euclidean norm.
  3. (3)
    If $v$ is a complex place, for any $P\in V(K_{v})$, the restriction to $E_{P}$ is the square of an hermitian norm.
  4. (4)
    There exists a finite number of places $S\subset M_{K}$ (which contains $M_{K}^{\infty}$) and a model $\mathcal{E}\rightarrow\mathcal{V}$ of $E\rightarrow V$ over $\mathcal{O}_{S}$ such that for any place $v\in M_{K}\setminus S$, for any $P\in\mathcal{V}(\mathcal{O}_{v})$,
    $\mathcal{E}_{P}=\{y\in E_{P}|\ \ ||y||_{v}\leqslant 1\}$
Remark 1.2.
For smooth, projective and split toric varieties over $\mathbf{Q}$ such model always exists over $\mathbf{Z}$.
The following example is particularly important, because, as we shall later see, the natural system of heights over toric varieties can be deduced from this example.
Example 1.3.
Let $N\in\mathbf{N}^{*}$. For $v\in M_{K}$, we consider the map:
$||-||_{v}:$ $K_{v}^{N+1}\rightarrow\mathbf{R}_{+}$
$(y_{0},..,y_{N})\longmapsto\max\limits_{i=0}^{N}|y_{i}|_{v}$
This defines an adelic norm $(||-||_{v})_{v\in M_{K}}$ over the line bundle $\mathcal{O}_{\mathbf{P}^{N}}(-1)\rightarrow\mathbf{P}^{N}$ which corresponds to the $\mathbf{G}_{m}$-torseur $\mathbf{A}^{N+1}\setminus\{0\}\rightarrow\mathbf{P}^{N}$.
Definition 1.4.
If $L\rightarrow V$ is a line bundle, $\{||-||_{v}\}$ an adelic norm over $L$, then we define the height $H_{L}:V(K)\rightarrow\mathbf{R}_{+}$ for any $x\in V(K)$ such that :
$H_{L}(x)=\prod\limits_{v\in M_{K}}||y||_{v}^{-1}$
for $y\in L^{\times}(K)$ over $x$.
Example 1.5.
In particular, by the previous example, for any $x\in K^{N+1}-\{0\}$, if we denote by $P=[x_{0}:..:x_{n}]\in\mathbf{P}^{N}(K)$ its image, we get:
$H_{\mathcal{O}_{\mathbf{P}^{N}}(1)}(P)=\prod\limits_{v\in M_{K}}\max\limits_{0\leq i\leq N}|x_{i}|_{v}.$
Now, we recall the definition of the group of Arakelov heights $\mathcal{H}(V)$ over $V$:
Definition 1.6.
Let $L$ and $L^{\prime}$ be two adelic line bundles over $V$ (i.e equipped with an adelic norm). Let $(||-||_{v})$ be the adelic norm over $L$.
We say that $L$ and $L^{\prime}$ are equivalent as adelic line bundles when there exists $M>0$ and a family $(\lambda_{v})_{v}$ of non negative real numbers such that:
  1. (1)
    $\{v|\lambda_{v}\neq 1\}$ is finite and $\prod\limits_{v}\lambda_{v}=1$
  2. (2)
    there exists an isomorphism of adelic line bundles from $L^{\otimes M}$ to $L^{\prime\otimes M}$, where the adelic norm on $L^{M}$ is induced by the adelic norm on $L$ given by $\{(\lambda_{v}.||-||_{v})_{v\in M_{K}}\}$.
We denote by $\mathcal{H}(V)$ the set of equivalence classes of adelic line bundles. This set is naturally equipped with a structure of group given by the tensorial product of adelic line bundle.
Hence, we can define the notion of system of heights and of multi-heights over $V$.
Definition 1.7.
A system of heights over $V$ corresponds to the data of a section $s:\operatorname{Pic}(V)\rightarrow\mathcal{H}(V)$ of the forgetful morphism
$\mathrm{o}:\mathcal{H}(V)\rightarrow\operatorname{Pic}(V)$
Such system of heights defines a multi-height map:
$h:V(K)\rightarrow\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$
such that for any $P\in V(K)$, $h(P)$ is the linear form defined by $L\in\operatorname{Pic}(V)$,
$h(P)(L)=\log H_{L}(P)$
In order to state the conjecture made by Peyre [17, question 4.8], we shall define a measure $\nu$ on $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$ as follows:
Definition 1.8.
[…]
for any a compact subset $\mathrm{D}$ of $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$, where the Haar measure $dy$ on $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$ is normalised so that the covolume of the dual lattice of the Picard group is one.
For any domain $\mathrm{D}\subset\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$, we write:
$V(K)_{h\in\mathrm{D}}=\{P\in V(K)\mid h(P)\in\mathrm{D}\}.$
Definition 1.9.
We shall say that $V$ is quasi-Fano if it verifies the following hypothesis:
  1. (1)
    A multiple of the class of $\omega_{V}^{-1}$ is the sum of an ample divisor and a divisor with normal crossings;
  2. (2)
    The set $V(\mathbf{Q})$ is Zariski dense;
  3. (3)
    The group $H^{i}(V,\mathcal{O}_{V})=\{0\}$ for $i=1,2$;
  4. (4)
    The geometric Brauer group $\operatorname{Br}(\overline{V})$ is trivial and the geometric Picard group $\operatorname{Pic}(\overline{V})$ has no torsion;
  5. (5)
    The closed cone $\text{C}_{\operatorname{eff}}(\overline{V})$ is generated by the classes of a finite set of effective divisors.
In particular, any Fano variety and any smooth, projective and split toric varieties are quasi-Fano. With these notations, Peyre states the following conjecture:
Conjecture 1.10.
We assume that our variety is quasi-Fano and that the effective cone $C_{\operatorname{eff}}(V)$ is generated by a finite number of line bundles. Let $\mathrm{D}_{1}$ be a compact polyhedron of $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$ and u be an element of the interior of the dual of the effective cone $(C_{\operatorname{eff}}(V)^{\vee})^{\circ}$. For a real number $B>1$, we set:
$\mathrm{D}_{B}=\mathrm{D}_{1}+\log(B)u$
There exists a ’small’ subset $T\subset V(K)$ so that we have an asymptotic behaviour of the form:
$\sharp(V(K)-T)_{h\in\mathrm{D}_{B}}\underset{B\rightarrow+\infty}{\sim}\nu(\mathrm{D}_{1})\beta(V)\tau(V)B^{\langle\omega_{V}^{-1},u\rangle}$
where $\beta(V)=\sharp H^{1}(\mathbf{Q},\operatorname{Pic}(V_{\overline{\mathbf{Q}}}))$ and $\tau(V)$ is the Tamagawa number of $V$.
The Tamagawa number of $V$, $\tau(V)$ is defined in 2.20.
1.2. The Manin-Peyre conjecture in the case of toric varieties
In this article all toric varieties are supposed to be smooth, projective and split toric variety over $\mathbf{Q}$. Let $X$ be such toric variety. We shall use in the rest of this paper that the anticanonical line bundle $\omega_{X}^{-1}$ is big, i.e is in the interior of the effective cone $\text{C}_{\text{eff}}(X)$.
Notations.
We denote by $T\subset X$ its open torus, by $\Sigma$ a fan associated to $X=X_{\Sigma}$, $\Sigma(1)$ the set of edges of $\Sigma$ and by $m$ the number of edges. We denote by $t$ the rank of the Picard group $\operatorname{Pic}(X)$. With our hypothesis, it is a free abelian group of finite rank.
To an edge $\rho\in\Sigma(1)$, we associate the $T$-invariant divisor $D_{\rho}$. Recall that
$\omega_{X}^{-1}=\sum\limits_{\rho\in\Sigma(1)}[D_{\rho}]$
We write
$\pi:\operatorname{Div}_{T}(X)=\mathbf{Z}^{\Sigma(1)}\rightarrow\operatorname{Pic}(X)$
the natural map which associates to a $T$-invariant divisor $D$ the class of the associated line bundle in $\operatorname{Pic}(X)$ which we write $[D]=\pi(D)$.
We denote by $N=X_{*}(T)$ the group of subgroups of one parameters of $T$ and by $M=X^{*}(T)$ its dual, the group of characters of $T$.
If $L$ is a line bundle over $X$, there exists a natural way to define an adelic norm $(||-||_{v})_{v\in M_{K}}$ on $L$, following [18, definition 9.2] (see also [3, proposition 2.1.2]). Let us explain how one can interpret this adelic norm in the case of $\mathbf{P}^{N}$.
Remark 1.11.
For $0\leq i\leq N$, we denote by $D_{i}$ the divisor $\mathbf{G}_{m}^{N}$-invariant of $\mathbf{P}^{N}$ corresponding to the hyperplane $\{X_{i}=0\}\subset\mathbf{P}^{N}$.
Then the adelic norm over $\mathcal{O}_{\mathbf{P}^{N}}(1)$ associated to $D_{i}$ corresponds to the one of the previous example 1.5.
Proof.
This is a consequence of [18, proposition 9.8]. ∎
Remark 1.12.
We can moreover observe that the construction of the adelic norm in [18, definition 9.2] over $\mathcal{O}(D)$ depends only on the class of the $T$-invariant divisor $D$ modulo $M=X^{*}(T)$, that is to say its class $[D]$ in $\operatorname{Pic}(X)$.
Proposition 1.13.
The heights system $s_{X}:\operatorname{Pic}(X)\rightarrow\mathcal{H}(X)$ defined on a toric variety $X$ is functorial for the morphisms of toric varieties.
[…]
${\mathcal{H}(Y)}$${\mathcal{H}(X)}$${\operatorname{Pic}(Y)}$${\operatorname{Pic}(X)}$$\scriptstyle{f^{*}}$$\scriptstyle{f^{*}}$$\scriptstyle{s_{Y}}$$\scriptstyle{s_{X}}$
We shall later see using versal torsor that this way of defining a height system over our toric variety $X$ (and a representative of its heights system) makes the count of rational points of our variety easier.
Notations.
Let $f$ be the smallest positive integer such that there exists $f$ rays of $\Sigma$ not contained in a cone of $\Sigma$. Note that $f\geqslant 2$.
The main theorem of this article states that the conjecture 1.10 for the multi-height asymptotic behaviour of rational points of toric varieties is true:
Theorem 1.14.
Let $\mathrm{D}_{1}$ be a finite union of compact polyhedrons in $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ and $u$ be an element of the interior of the dual of the effective cone $(C_{\operatorname{eff}}(X)^{\vee})^{\circ}$. For a real number $B>1$, we set:
$\mathrm{D}_{B}=\mathrm{D}_{1}+\log(B)u.$
Let $\varepsilon\in]0,1-\frac{1}{f}[$. Then we have an asymptotic behaviour of the form:
$\sharp(X(\mathbf{Q}))_{h\in\mathrm{D}_{B}}\underset{B\rightarrow+\infty}{=}\nu(\mathrm{D}_{1}).\tau(X).B^{\langle\omega_{X}^{-1},u\rangle}\left(1+O\left(B^{-(1-\frac{1}{f}-\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}\right)\right)$
where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 1.8.
Example 1.15.
If $\mathbf{b}=(L_{1},..,L_{r})$ is a family of line bundles such that there classes $\mathbf{R}$-generate the Picard group $\operatorname{Pic}(X)_{\mathbf{R}}$, $\alpha_{1},..,\alpha_{r},\beta_{1},..,\beta_{r}$ are real numbers such that for any $i$, $\alpha_{i}<\beta_{i}$, then we define a compact polyhedron of $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$:
$\mathrm{D}_{1}=\{a\in\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}\mid\alpha_{i}\leqslant a(L_{i})\leqslant\beta_{i}\}.$
All compact polyhedrons of $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ are of this form.
2. The notion of universal torsors
In this section, we shall present how the study of the multi-height distribution of rational points of toric varieties naturally lifts to a counting problem in the universal torsor at the origin.
2.1. Generalities about versal torsors
Originally, Colliot-Thélène and Sansuc defined universal torsors using a cohomological characterization (see [6, section 2]) that can be generalized over an noetherian base scheme $S$. For this paragraph, we essentially refer to [18, section 5].
We suppose that $f:V\rightarrow S$ is a smooth proper surjective morphism of constant relative dimension with geometrically connected fibers such that $R^{2}f_{*}\mathcal{O}_{X}=0$ and the relative Picard functor $\operatorname{Pic}_{V/S}$ is étale locally the constant sheaf associated with an ordinary free abelian group of finite rank. We also suppose that $f$ has a rational section. We shall denote by $T_{\operatorname{NS}}$ the group scheme associated to $\operatorname{Pic}_{V/S}$ by duality. Then we have the following exact sequence, functorial in $S$ and $V$ (see [18, section 5, p.163] which generalize [6, 2.0.2]):
[…]
Definition 2.1.
A $T_{\operatorname{NS}}$-torsor $\mathcal{T}\rightarrow V$ is said to be universal if its class
$[\mathcal{T}]\in H^{1}(V,T_{\operatorname{NS}})$
is sent to $\operatorname{Id}_{\operatorname{Pic}_{V/S}}$ via $\chi$.
[…]
$\chi([\mathcal{T}]):$ $H^{1}(V,\mathbf{G}_{m})\rightarrow H^{1}(V,G_{m})$
$[L^{\times}]\longmapsto[\mathcal{T}\overset{T_{\operatorname{NS}}}{\times_{\phi_{L}}}\mathbf{G}_{m}]$
where $\phi_{L}$ is the character of $T_{\operatorname{NS}}$ associated to $[L^{\times}]\in H^{1}(V,\mathbf{G}_{m})$ i.e $[L]\in\operatorname{Pic}(V)$.
Let us choose $\mathbf{b}=(L_{1},..,L_{t})$ a basis of $\operatorname{Pic}(V)$ and write
$\mathcal{T}_{\mathbf{b}}=L_{1}^{\times}\times_{V}...\times_{V}L_{t}^{\times}$
The torus $T_{\operatorname{NS}}$ is isomorphic to $\mathbf{G}_{m}^{t}$ via the characters given by the basis $\mathbf{b}$. Hence we get a natural left action of $T_{\operatorname{NS}}$ over
$q:\mathcal{T}_{\mathbf{b}}\rightarrow V$
which makes $\mathcal{T}_{\mathbf{b}}$ a $T_{\operatorname{NS}}$-torsor over $V$.
In particular, provided that we have chosen the right convention for the contracted product, we find that $\chi([\mathcal{T}_{\mathbf{b}}])=\operatorname{Id}_{\operatorname{Pic}_{V/S}}$ (otherwise it will be $-\operatorname{Id}_{\operatorname{Pic}_{V/S}}$).
By [16, proposition 8.1], we get that the previous construction is independent, up to a $T_{\operatorname{NS}}$-equivariant isomorphism, to the choice of the basis of $\operatorname{Pic}(V)$ chosen.
Remark 2.2.
By functoriality of this cohomological characterization, if $\mathcal{T}\rightarrow V$ is a universal torsor and $S^{\prime}\rightarrow S$ is a base change, then $\mathcal{T}_{S^{\prime}}\rightarrow V_{S^{\prime}}$ is again a universal torsor.
Let us recall the following proposition from [18, proposition 5.15], which is valid under more general assumptions that our hypothesis that $\operatorname{Pic}_{V/S}$ is a constant sheaf:
Proposition 2.3.
There is a universal torsor over $V$. Moreover the isomorphism classes of universal torsors over $V$ are parameterized by the $H^{1}(S,T_{\operatorname{NS}})$-orbits in $H^{1}(V,T_{\operatorname{NS}})$ defined by $\chi^{-1}(\operatorname{Id})$.
Cohomological arguments give that if $\alpha\in H^{1}(V,T_{\operatorname{NS}})$ is the class of a universal torsor $\mathcal{T}\rightarrow V$, $x\in V(S)$ and $P\rightarrow S$ is a $T_{\operatorname{NS}}$-torsor associated to the class $\alpha(x)$ then $\mathcal{T}\overset{T_{\operatorname{NS}}}{\times}P\rightarrow V$ is a universal torsor such that the fiber over $x$ has a rational point. This come from the fact that if $\beta\in H^{1}(V,T_{\operatorname{NS}})$ is the class of $\mathcal{T}\overset{T_{\operatorname{NS}}}{\times}P\rightarrow V$, we have by construction $\beta(x)=0$.
Suppose again that $S=\operatorname{Spec}(\mathbf{Z})$ and that $\operatorname{Pic}_{V/\mathbf{Z}}$ is the constant sheaf associated to a free abelian group of finite rank $t$. As $\mathbf{Z}$ is a principal ring, we have that $H^{1}(\mathbf{Z},T_{\operatorname{NS}})=\{1\}$, there is up to isomorphism a universal torsor $q:\mathcal{T}\rightarrow V$. Moreover, we have:
$q(\mathcal{T}(\mathbf{Z}))=V(\mathbf{Z}).$
2.2. Local heights defined over the universal torsors
We keep the hypothesis of the previous paragraph on $V$. We write $q:\mathcal{T}\rightarrow V$ for the universal torsor we have chosen (unique up to isomorphism by hypothesis). We fix $y_{0}\in\mathcal{T}(\mathbf{Q})$ and we denote by $x_{0}\in V(\mathbf{Q})$ its image via $q$.
We know ([17, remark 4.19]) that the universal torsor over $V_{\mathbf{Q}}$ is then characterized by a universal property as a pointed torsor.
Proposition 2.4.
For any pointed torsor $P\rightarrow V$ which send $p_{0}\in P(\mathbf{Q})$ to $x_{0}\in V(\mathbf{Q})$, under a multiplicative group $T$, there exists a unique group morphism $\varphi:T_{NS}\rightarrow T$ and a unique $\varphi$équivariant morphism $\Phi:\mathcal{T}\rightarrow P$ over $V$ which is compatible with the marqued point. That is to say the following diagram is commutative:
${{(\mathcal{T},y_{0})}}$${{(P,p_{0})}}$${{(V,x_{0})}}$$\scriptstyle{\exists!\Phi}$
We recall the construction of Peyre [17, construction 4.27] to define the local height associated to the universal torsor. Suppose we have a given heights system
$s:\operatorname{Pic}(V)\rightarrow\mathcal{H}(V)$
We fix a place $v_{0}\in M_{\mathbf{Q}}$. For any linde bundle $L$, we have by the previous proposition a unique morphism $\Phi_{L}:\mathcal{T}\rightarrow L^{\times}$ which is equivariant for the group morphism $\chi_{L}:T_{NS}\rightarrow\mathbf{G}_{m}$ given by $[L]\in X^{*}(T_{NS})=\operatorname{Pic}(V)$. This morphism is unique up to multiplication by a scalar $\lambda\in\mathbf{Q}^{\times}$.
For each $[L]$ in $\operatorname{Pic}(V)$, we choose a representative $(||-||_{v})$ of the adelic norm associated to $s([L])$. Then we define for any $v\in M_{\mathbf{Q}}$ the map:
$||-||_{L,v}:\mathcal{T}$ $(\mathbf{Q}_{v})\rightarrow\mathbf{R}$
$y\longmapsto\frac{||\phi_{L}(y)||_{v}}{||\phi_{L}(y_{0})||_{v}}\text{ si }v\neq v_{0}$
$y\longmapsto\frac{||\phi_{L}(y)||_{v_{0}}}{||\phi_{L}(y_{0})||_{v_{0}}}H(L,q(y_{0}))^{-1}$
Definition 2.5.
The local height in the place $v\in M_{\mathbf{Q}}$ associated to the universal torsor $\mathcal{T}$ is defined on $\mathcal{T}(\mathbf{Q}_{v})$ by:
$H_{\mathcal{T},v}(y,L)=||y||_{L,v}^{-1}$
We also define a local logarithm height:
$h_{\mathcal{T},v}:\mathcal{T}(\mathbf{Q}_{v})\rightarrow\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$
such that for any $L\in\operatorname{Pic}(V)$, we have:
$H_{\mathcal{T},v}(y,L)=q_{v}^{-\langle h_{\mathcal{T},v}(y),L\rangle}.$
In particular, we have for any $L\in\operatorname{Pic}(V)$ and for any $y\in\mathcal{T}(\mathbf{Q})$,
(2.1) $H_{L}(q(y))=\prod\limits_{v\in M_{\mathbf{Q}}}H_{\mathcal{T},v}(y,L).$
In order to describe one of the key property of this local logarithm height, we need to recall the definition of the local logarithm map associated to a torus.
Definition 2.6.
Let $T$ be a torus over $\mathbf{Q}$, we denote $M=X^{*}(T)$ the $\operatorname{Gal}(\bar{\mathbf{Q}}/\mathbf{Q})$-lattice corresponding to the character group of $T$ and by $N=X_{*}(T)$ its dual for the usual pairing. For $v\in M_{\mathbf{Q}}$, we have a natural map:
$\log_{T,v}:T({\mathbf{Q}}_{v})\rightarrow N_{\mathbf{R}}$
such that for any $\chi\in M$, for any $t\in T({\mathbf{Q}}_{v})$, we have:
$|\chi(t)|_{v}=q_{v}^{\langle\log_{T,v}(t),\chi\rangle}.$
We can then write the following lemma ([17, lemme 4.30]):
Lemma 2.7.
For any $t\in T_{\operatorname{NS}}({\mathbf{Q}}_{v})$, for any $y\in\mathcal{T}({\mathbf{Q}}_{v})$,
$h_{\mathcal{T},v}(t.y)=\log_{T_{NS},v}(t)+h_{\mathcal{T},v}(y).$
2.3. Universal torsors for split, smooth and projective toric varieties
We keep the notations from Notations, let $X$ be a toric variety over $\mathbf{Q}$ (smooth, projective and split) and denote by $\Sigma$ an associated fan. In our particular case, the universal torsor $\mathcal{T}_{\Sigma}\rightarrow X$ can be described as an open subset of $\mathbf{A}^{\Sigma(1)}$ obtained by cutting out a certain closed subset $Z(\Sigma)$.
As in [7, Chapter 5, §1], we define for a cone of the fan $\sigma\in\Sigma$, the monomial:
$X^{\tilde{\sigma}}=\prod\limits_{\rho\not\in\sigma(1)}X_{\rho}.$
Then we shall use the following definition of the exceptional set $Z(\Sigma)$:
Definition 2.8.
We denote by $B(\Sigma)$ the ideal of $\mathbf{Z}[X_{\rho},\rho\in\Sigma(1)]$ generated by the monomials $X^{\tilde{\sigma}}=\prod\limits_{\rho\not\in\sigma(1)}X_{\rho}$ where $\sigma$ belongs to $\Sigma_{\max}$ the set of cones of the fan $\Sigma$ of maximal dimension. Then we define the closed subset of $\mathbf{A}^{\Sigma(1)}$ as:
$Z(\Sigma)=V(B(\Sigma))=\bigcap\limits_{\sigma\in\Sigma_{\max}}\left(\bigcup\limits_{\rho\not\in\sigma(1)}\left(X_{\rho}=0\right)\right).$
We define $\mathcal{T}_{\Sigma}=\mathbf{A}^{\Sigma(1)}-Z(\Sigma)$.
In particular, this means that for $y\in\mathbf{Z}^{\Sigma(1)}-\{0\}$, we have the following equivalence:
$y\in\mathcal{T}_{\Sigma}(\mathbf{Z})\Leftrightarrow\gcd(y^{\tilde{\sigma}},\sigma\in\Sigma_{\max})=1.$
We shall give here an other description of the universal torsor $\mathcal{T}_{\Sigma}\rightarrow X$ that we shall not use in this paper but which is more intuitive. It is an open subset of $\mathbf{A}^{\Sigma(1)}$ obtained by cutting out a certain closed subset which can be described in term of primitive collections which are defined as follows (see [7, definition 5.1.5]):
Definition 2.9.
A subset $C\subset\Sigma(1)$ is a primitive collection if:
  • $C\not\subset\sigma(1)$ for any $\sigma\in\Sigma$
  • for every proper subset $C^{\prime}\subsetneq C$, there exists $\sigma\in\Sigma$ such that $C^{\prime}\subset\sigma(1)$
Hence $Z(\Sigma)$ can be described as a union of irreducible components as in [7, proposition 5.1.6]:
Proposition 2.10.
$Z(\Sigma)=\bigcup\limits_{C}V(x_{\rho}\mid\rho\in C)$
where $C$ belongs to the set of primitive collections.
Let $\beta:\operatorname{Div}_{T}(X)^{\vee}=\mathbf{Z}^{\Sigma(1)}\rightarrow N$ be the $\mathbf{Z}$-linear map which sends the vector of the canonical basis $e_{\rho}$ to $u_{\rho}$ the minimal generator of the edge $\rho$. Denote by $\Sigma_{0}$ the fan of $\mathcal{T}_{\Sigma}$ (see [7, proposition 5.19]). Both toric varieties $\mathcal{T}_{\Sigma}$ and $X$ are defined over $\mathbf{Z}$ via their fans. We state here the following essential proposition:
Proposition 2.11.
The natural toric morphism $q:\mathcal{T}_{\Sigma}\rightarrow X$ over $\mathbf{Z}$ given by $\beta:(\mathbf{Z}^{\Sigma(1)},\Sigma_{0})\rightarrow(N,\Sigma)$ makes $\mathcal{T}_{\Sigma}$ the universal torsor (unique up to isomorphism) over $X$.
Proof.
See [18, lemma 5.2 and section 10]. ∎
Notations.
We shall denote by $T_{\Sigma(1)}$ the open torus of the toric variety $\mathcal{T}_{\Sigma}$, which is naturally isomorphic to $\mathbf{G}_{m}^{\Sigma(1)}$ via the character
$\chi_{D_{\rho}}:$ $T_{\Sigma(1)}\rightarrow\mathbf{G}_{m}$
$y\longmapsto y_{\rho}.$
Note that we have $q^{-1}(T)=T_{\Sigma(1)}$.
2.3.1. An alternative description of the system of heights over toric varieties
Recall we have the following exact sequence associated to the toric variety $X$:
[…]
To an effective divisor $D=\sum\limits_{\rho\in\Sigma(1)}a_{\rho}D_{\rho}$, we associate the global section $X^{D}$ in the Cox Ring $\mathbf{Q}[X_{\rho}|\rho\in\Sigma(1)]$ of the toric variety $X_{\Sigma}$ defined by:
$X^{D}=\prod\limits_{\rho\in\Sigma(1)}X_{\rho}^{a_{\rho}}.$
Using Kleiman’s criterion (see [9, theorem 1.27]), we can choose $(L_{1},..,L_{t})$ a basis of $\operatorname{Pic}(X)$ where each $L_{i}$ is an ample line bundle. Moreover by [7, theorem 6.1.16], because $X$ is smooth, $L_{i}$ ample implies $L_{i}$ very ample. We get the following lemma:
Lemma 2.12.
There exists a basis $L_{1},..,L_{t}$ of $\operatorname{Pic}(X)$ such that each $L_{i}$ is a very ample line bundle.
We choose such basis. For $1\leqslant k\leqslant t$, we set $h^{0}(X,L_{k})=r_{k}$ and we take $\mathcal{M}_{k}\subset\operatorname{Div}_{T}(X)$ a family of $T$-invariant divisor such that:
(2.2) $H^{0}(X,L_{k})=\bigoplus\limits_{D\in\mathcal{M}_{k}}\mathbf{Q}.X^{D}.$
The fact that such family exists is a consequence of the description of the Cox Ring of the toric variety $X$. In fact, one can show that $\mathcal{M}_{k}=\pi^{-1}(L_{k})\cap\mathbf{N}^{\Sigma(1)}$.
We have the following lemma:
Lemma 2.13.
The closed immersion given by the very ample line bundles:
$X_{\mathbf{Q}}\rightarrow\prod\limits_{i=1}^{t}\mathbf{P}^{r_{i}-1}_{\mathbf{Q}}$
is in fact defined over $\mathbf{Z}$ and we have the following commutative diagram which is cartesian:
${\mathcal{T}_{\mathbf{b}}}$${\prod\limits_{i=1}^{t}\mathbf{A}^{r_{i}}\setminus\{0\}}$${X}$${\prod\limits_{i=1}^{t}\mathbf{P}^{r_{i}-1}}$
Let us denote by $A_{k}\in\mathcal{M}_{r_{k},m}(\mathbf{Z})$ the matrix whose lines are given by the coordinates of $D\in\mathcal{M}_{k}$ in the basis $\{D_{\rho}|\rho\in\Sigma(1)\}$.
Proposition 2.14.
The $\mathbf{Z}$-linear map given by the matrix $\Bigg{[}\begin{smallmatrix}&A_{1}\\ &\vdots\\ &A_{t}\end{smallmatrix}\Bigg{]}$, that is to say:
$\Phi:\mathbf{Z}^{\Sigma(1)}\rightarrow\bigoplus\limits_{1\leqslant i\leqslant t}\mathbf{Z}^{r_{i}}$
induced the toric morphism of toric varieties:
$\mathcal{T}_{\Sigma}\rightarrow\prod\limits_{i=1}^{t}\mathbf{A}^{r_{i}}\setminus\{0\}$
which corresponds to the cartesian diagram 2.13. Moreover, the induced morphism by quotient
$X\rightarrow\prod\limits_{i=1}^{t}\mathbf{P}^{r_{i}-1}$
is again a toric morphism of toric varieties.
Hence using the height system previously defined for toric varieties, fixing the point $1=(1)_{\rho\in\Sigma(1)}$ of $\mathcal{T}_{\Sigma}(\mathbf{Q})$, we can describe the local height at the universal torsor $\mathcal{T}_{\Sigma}$. We use the functorial property 1.13 of the notion of heights system for toric morphism and we get that for any $v\in M_{\mathbf{Q}}$ and any $k\in\{1,..,t\}$ :
$H_{\mathcal{T},v}(y,L_{k})=\max\limits_{m\in\mathcal{M}_{k}}|X^{m}(y)|.$
In fact using the description of the universal torsor given by the cartesian diagram 2.13, we can deduce an other description of the $\mathbf{Z}$-points of $\mathcal{T}_{\Sigma}$:
Theorem 2.15.
$\mathcal{T}_{\Sigma}(\mathbf{Z})=\{y\in\mathbf{Z}^{\Sigma(1)}|\forall\ i\in\{1,..,t\},\ \gcd\limits_{D\in\mathcal{M}_{i}}(y^{D})=1\}$
In particular, for any $y\in\mathcal{T}_{\Sigma}(\mathbf{Z})$, we have for $1\leqslant i\leqslant t$,
$H_{L_{i}}(q(y))=\max\limits_{D\in\mathcal{M}_{i}}(|y^{D}|)=H_{\mathcal{T},\infty}(y,L_{i}).$
As $L_{1},..,L_{t}$ form a basis of $\operatorname{Pic}(X)$, we get also the following corollary:
Corollary 2.16.
For any $y\in\mathcal{T}_{\Sigma}(\mathbf{Z})$, we have:
$h(q(y))=h_{\mathcal{T},\infty}(y).$
2.3.2. The action of the Neron-Severi torus on each coordinate
In this paragraph, we shall describe the action of $T_{\operatorname{NS}}$ on $\mathcal{T}_{\Sigma}$. It is induced by the group morphism $\pi^{*}:T_{\operatorname{NS}}\rightarrow T_{\Sigma(1)}$, dual to the morphism $\pi:\operatorname{Div}_{T}(X)\rightarrow\operatorname{Pic}(X)$.
Recall that we have a group morphism
$\exp:\operatorname{Pic}(X)$ ${}^{\vee}_{\mathbf{R}}\hookrightarrow\mathbf{T}_{\operatorname{NS}}(\mathbf{R})$
$a\longmapsto e^{a}$
such that for any $L\in\operatorname{Pic}(X)$, $\chi_{L}(e^{a})=e^{a(L)}$ where $\chi_{L}\in X^{*}(T_{\operatorname{NS}})$ is the character associated to $L\in\operatorname{Pic}(X)$.
To any $T$-invariant divisor $D$, we can associate a character $\chi_{D}:T_{\Sigma(1)}\rightarrow\mathbf{G}_{m}$. The composition
$\chi_{D}\circ\pi^{*}:T_{\operatorname{NS}}\rightarrow\mathbf{G}_{m}$
corresponds to the character $\chi_{[D]}$ where $[D]\in\operatorname{Pic}(X)$ is the class of $D\in\operatorname{Div}_{T}(X)$.
For any scheme $S$, for any $\rho\in\Sigma(1)$, the action of $t\in T_{\operatorname{NS}}(S)$ on the coordinate of $\mathbf{A}^{\Sigma(1)}$ corresponding to the edge $\rho$ is given by multiplication by $\chi_{[D_{\rho}]}(t)$ as $\chi_{D_{\rho}}\circ\pi^{*}=\chi_{[D_{\rho}]}$. Then we can deduce the following proposition:
Proposition 2.17.
Let $a\in\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$. Then for any $\rho\in\Sigma(1)$, the action of $e^{a}\in T_{\operatorname{NS}}(\mathbf{R})$ on the coordinate of $\mathbf{A}^{\Sigma(1)}$ corresponding to the edge $\rho$ is given by multiplication by $e^{\langle[D_{\rho}],a\rangle}$.
2.3.3. A useful covering of the real variety associated to our toric variety
In this paragraph, we suppose that $X(\mathbf{R})$ is equipped with the real topology. Then we shall show the following theorem:
Theorem 2.18.
The restriction map given by $h_{\mathcal{T},\infty}^{-1}(\{0\})\rightarrow X(\mathbf{R})$ is a finite covering.
Proof.
Let $P\in X(\mathbf{R})$ and let $U$ be an open subscheme of $X_{\mathbf{R}}$ such that $P\in U(\mathbf{R})$ and there exists a section $s:U\rightarrow\mathcal{T}_{\mathbf{R}}$ of $q:\mathcal{T}_{\mathbf{R}}\rightarrow X_{\mathbf{R}}$. Such section exists because by Hilbert’s 90, we have $H^{1}_{\text{et}}(X_{\mathbf{R}},T_{\operatorname{NS}})=H^{1}_{\text{Zar}}(X_{\mathbf{R}},T_{\operatorname{NS}})$. Hence we have a continuous section
$P\in U(\mathbf{R})\longmapsto e^{-h_{\mathcal{T},\infty}(s(P))}.s(P)$
[…]
$U(\mathbf{R})$$\times\ker(\log_{T_{\operatorname{NS}},\infty})\rightarrow p^{-1}(U(\mathbf{R}))$
$(P,t)\longmapsto t.e^{-h_{\mathcal{T},\infty}(s(P))}.s(P).$
As $\ker(\log_{T_{\operatorname{NS}},\infty})\simeq\{-1,1\}^{t}$, it is a finite covering. ∎
Corollary 2.19.
$h_{\mathcal{T},\infty}^{-1}(\{0\})$ is compact.
Proof.
By [5, Chapitre I, §4, n°5, théorème 1], the map $h_{\mathcal{T},\infty}^{-1}(\{0\})\rightarrow X(\mathbf{R})$ is proper. Moreover, $X(\mathbf{R})$ is compact for the real topology as it is a projective variety. Hence we get our corollary. ∎
2.4. Measure over the universal torsors and Tamagawa number of our variety
2.4.1. Adelic measure and Tamagawa number over a variety
Here $V$ means a smooth, separated scheme which have geometrically connected fibers over $\operatorname{Spec}(\mathbf{Z})$ and such that $V_{\mathbf{Q}}$ is a smooth, projective and geometrically integral variety over $\mathbf{Q}$. Let us recall here how an adelic metric $(||-||_{v})_{v\in M_{\mathbf{Q}}}$ over $\omega_{V}^{-1}$ defines a measure $\mu_{V}$ over
$V(A_{\mathbf{Q}})=\prod\limits_{v\in M_{\mathbf{Q}}}V(\mathbf{Q}_{v}).$
As in [13, 2.2.1], we can define locally a measure over $V(\mathbf{Q}_{v})$ by the relation
$\|\frac{\partial}{\partial x_{1,v}}\wedge..\wedge\frac{\partial}{\partial x_{n,v}}\|_{v}dx_{1,v}..dx_{n,v}$
where $(x_{1},..,x_{n}):\Omega\rightarrow\mathbf{Q}_{v}^{n}$ is a local system of coordinates defined on an open subset $\Omega\subset V(\mathbf{Q}_{v})$ and $\frac{\partial}{\partial x_{1,v}}\wedge..\wedge\frac{\partial}{\partial x_{n,v}}$ is seen as a local section of $\omega_{V}^{-1}$. This local measure does not depend on the choice of coordinates, therefore we can glue together these measures and we get a measure $\mu_{V,v}$ over $V(\mathbf{Q}_{v})$.
By [13, lemma 2.1.2], we have that if $V$ has good reduction at every finite places and the norm is defined by the model (for example if it is a split toric variety), then:
$\mu_{V,v}(V(\mathbf{Q}_{v}))=\frac{\sharp V(\mathbf{F}_{v})}{p_{v}^{\dim(V)}}.$
In order to define the Tamagawa number of our variety $V$, one adds some convergence factors $\lambda_{v}=L_{v}(1,\operatorname{Pic}(\overline{V}))$ for $v$ a finite place and $\lambda_{v}=1$ if $v$ is infinite.
Definition 2.20.
The Tamagawa number of $V$ is here defined as $\tau(V)=\mu_{V}(V(A_{\mathbf{Q}}))$ where $\mu_{V}=\prod\limits_{v\in M_{\mathbf{Q}}}\lambda_{v}^{-1}\mu_{V,v}$.
Remark 2.21.
In the case where $V$ is split, we have that for any $v\in M_{\mathbf{Q}}^{0}$,
$\lambda_{v}=(1-\frac{1}{p})^{t}.$
We get that if moreover $V$ has a good reduction at every finite places:
$\tau(V)=\mu_{V,\infty}(V(\mathbf{R}))\prod\limits_{p\in\mathbb{P}}(1-\frac{1}{p})^{t}\frac{\sharp V(\mathbf{F}_{p})}{p^{\dim(V)}}.$
Before defining a natural measure over universal torsor, let us recall as in [14, definition 3.1.2], that one can define a natural measure over a split torus $T$:
Definition 2.22.
Let $(\xi_{1},..,\xi_{t})$ be a basis of $X^{*}(T)_{\mathbf{R}}$ then the differential form
$\omega_{T}=\bigwedge\limits_{i=1}^{t}\frac{d\xi_{i}}{\xi_{i}}$
does not depend on the choice of the basis, up to its sign. It is called the canonical form of $T$. This form defines a natural measure over $T(\mathbf{R})$.
2.4.2. Adelic measure over universal torsors
In order to make the notations less cluttered, for the rest of this article, we shall write $\mathcal{T}=\mathcal{T}_{\Sigma}$ where $\mathcal{T}_{\Sigma}$ is the universal torsor defined in 2.11 over the toric variety $X$. In this paragraph, we recall the construction from [17, 4.3.2] to define for any $v\in M_{\mathbf{Q}}$, a measure $\omega_{\mathcal{T},v}$ over $\mathcal{T}(\mathbf{Q}_{v})$.
Using the exact sequence
$1\rightarrow\mathbf{G}_{m,X}\rightarrow q_{*}G_{m,\mathcal{T}}\rightarrow\widehat{T_{\operatorname{NS}}}\rightarrow 0$
associated to the universal torsor, which corresponds up to sign to the element $[\mathcal{T}]\in H^{1}(X,T_{\operatorname{NS}})$ seen as an element in $\operatorname{Ext}^{1}_{X}(\mathbf{G}_{m,X},T_{\operatorname{NS}})$ (see [6, 1.4.3]), we get that $H^{0}(\mathcal{T},\mathbf{G}_{m})=\mathbf{Q}^{\times}$ and that $\operatorname{Pic}(\mathcal{T})$ is trivial. Moreover, by [15, lemme 2.1.10], $\omega_{\mathcal{T}}$ is isomorphic to the pull-back of $\omega_{X}$. Hence there exists a global section of $\omega_{\mathcal{T}}$ corresponding to a trivialization of the line bundle so we get the following proposition [17, proposition 4.24]:
Proposition 2.23.
Up to multiplication by a scalar, there exists a unique volume form over $\mathcal{T}$.
We follow [17, construction 4.25] in order to describe locally the measure $\omega_{\mathcal{T},v}$ over $\mathcal{T}(\mathbf{Q}_{v})$. Let us choose a non zero section $\omega$ of $\omega_{\mathcal{T}}$. For any place $v\in M_{\mathbf{Q}}$, the expression
$\left|\left\langle\omega,\frac{\partial}{\partial x_{1}}\wedge..\wedge\frac{\partial}{\partial x_{m}}\right\rangle\right|_{v}dx_{1,v}..dx_{m,v}$
defined a local measure. Gluing this local measures, we get a measure $\omega_{\mathcal{T},v}$ over $\mathcal{T}(\mathbf{Q}_{v})$.
Remark 2.24.
In our case, recall that $X$ is a smooth, projective and split toric variety and $\mathcal{T}$ an open subset of $\mathbf{A}^{\Sigma(1)}$, the previous construction gives us that $\omega_{\mathcal{T},v}$ is the restriction of the normalised Haar measure $dx_{v}$ over $\mathbf{Q}_{v}^{\Sigma(1)}$ to the open subset $\mathcal{T}(\mathbf{Q}_{v})$ if $v\in M_{\mathbf{Q}}^{0}$ and is the restriction of the natural Lesbesgue measure $dx_{\infty}$ over $\mathbf{R}^{\Sigma(1)}$ to $\mathcal{T}(\mathbf{R})$.
Then using the basic property of the universal torsor, we have the following theorem:
Theorem 2.25.
For any $v\in M_{\mathbf{Q}}^{0}$, we have:
$\omega_{\mathcal{T},v}(\mathcal{T}(\mathbf{Z}_{v}))=\frac{\sharp\mathcal{T}(\mathbf{F}_{v})}{p_{v}^{\dim(\mathcal{T})}}=(1-p_{v})^{t}.\frac{\sharp X(\mathbf{F}_{v})}{p_{v}^{\dim(X)}}.$
To end this section, we shall compute $\omega_{\mathcal{T},\infty}(\mathcal{D})=\operatorname{Vol}(\mathcal{D})$ where
$\mathcal{D}=\{y\in\mathcal{T}(\mathbf{R})\mid h_{\mathcal{T},\infty}(y)\in\mathrm{D}\}$
for a compact subset $\mathrm{D}\subset\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$. We shall connect it to $\nu(\mathrm{D})$ where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 1.8. We shall need the following lemma which was stated in [14, 4.4.1]:
Lemma 2.26.
Let $P\mapsto y_{P}$ be any measurable section of $\mathcal{T}(\mathbf{R})\rightarrow X(\mathbf{R})$. For any function $f:X(\mathbf{R})\rightarrow\mathbf{R}$ compactly supported and measurable, the function:
$P\longmapsto\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt)$
does not depend on the choice of the section $P\mapsto y_{P}$ and we have:
$\int\limits_{\mathcal{T}(\mathbf{R})}f(y)\omega_{\mathcal{T},\infty}(dy)=\int\limits_{X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt).$
Proof.
The first assertion is a consequence of the invariance of the Haar measure (see [4, Chapitre VII, §2, n°2, proposition 1]). To show the second assertion, one must remark that in [18], Salberger shows that if you define the measure
$\omega_{\text{Sal},\infty}(dy)=\frac{\omega_{\mathcal{T},\infty}(dy)}{H_{\mathcal{T},\infty}(y,\omega_{X}^{-1})},$
it is a $T_{\operatorname{NS}}(\mathbf{R})$-invariant measure and the quotient measure $\frac{\omega_{\text{Sal},\infty}}{\omega_{T_{\operatorname{NS}}}}$ is $\mu_{X,\infty}$. Hence by [4, Chapitre VII, §2, n°2, proposition 4], we have that for any $f:X(\mathbf{R})\rightarrow\mathbf{R}$ compactly supported function, we may choose a section $P\mapsto y_{P}$ such that
$H_{\mathcal{T},\infty}(y_{P},\omega_{X}^{-1})=1$
for any $P$, and we get:
$\int\limits_{\mathcal{T}(\mathbf{R})}f(y)\omega_{\mathcal{T},\infty}(dy)$ $=\int\limits_{\mathcal{T}(\mathbf{R})}f(y)H_{\mathcal{T},\infty}(y,\omega_{X}^{-1})\omega_{\text{Sal},\infty}(dy)$
$=\int\limits_{X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}H_{\mathcal{T},\infty}(t.y_{P},\omega_{X}^{-1})f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt)$
$=\int\limits_{X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}f(t.y_{P})\omega_{T_{\operatorname{NS}}}(dt).$
Proposition 2.27.
$\omega_{\mathcal{T},\infty}(\mathcal{D})=\operatorname{Vol}(\mathcal{D})=2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}).$
Proof.
Applying the previous lemma for $f=\mathbf{1}_{\mathcal{D}}$, we get:
$\omega_{\mathcal{T},\infty}(\mathcal{D})=\int\limits_{P\in X(\mathbf{R})}\mu_{X,\infty}(dP)\int\limits_{T_{\operatorname{NS}}(\mathbf{R})}|\chi_{\omega_{X}^{-1}}(t)|_{\infty}\mathbf{1}_{\log_{T_{\operatorname{NS}},\infty}^{-1}(\mathrm{D})}(t)\omega_{T_{\operatorname{NS}}}(t)$
where for each $P\in X(\mathbf{R})$, we have choosen $y_{P}\in\mathcal{T}_{P}(\mathbf{R})$ such that $h_{\mathcal{T},\infty}(y_{P})=0$. Then using that we have the following exact sequence of locally compact groups:
$0\rightarrow(\{\pm 1\}^{t},\delta_{\text{compt}})\rightarrow(T_{\operatorname{NS}}(\mathbf{R}),\omega_{T_{\operatorname{NS}}})\xrightarrow{\log_{T_{\operatorname{NS}},\infty}}(\operatorname{Pic}(X)^{\vee}_{\mathbf{R}},da)\rightarrow 0$
where $da$ is the Lebesgue measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$, we get:
[…]
$=2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}).$
2.5. The Möbius function associated to the universal torsor
For this section, we follow the notations of Salberger [18, Notations 11.8] in order to introduce the Möbius function associated to the universal torsor $\mathcal{T}$ over $X$.
Notations.
  1. (1)
    $\mathbf{1}_{\mathcal{T}(\mathbf{Z})}:(\mathbf{N}^{*})^{\Sigma(1)}\rightarrow\{0,1\}$ is the indicator function of $\mathcal{T}(\mathbf{Z})\cap(\mathbf{N}^{*})^{\Sigma(1)}$, that is to say we have the following equivalence:
    $\mathbf{1}_{\mathcal{T}(\mathbf{Z})}(e)=1\Leftrightarrow\gcd(y^{\tilde{\sigma}},\sigma\in\Sigma_{\max})=1.$
  2. (2)
    for $d\in(\mathbf{N}^{*})^{\Sigma(1)}$, we write $\mathbf{1}_{d.\mathbf{Z}^{\Sigma(1)}}:(\mathbf{N}^{*})^{\Sigma(1)}\rightarrow\{0,1\}$ the indicator function of $\bigoplus\limits_{\rho\in\Sigma(1)}d_{\rho}.\mathbf{Z}\cap(\mathbf{N}^{*})^{\Sigma(1)}$, that is to say we have the following equivalence:
    $\mathbf{1}_{d.\mathbf{Z}^{\Sigma(1)}}(e)=1\Leftrightarrow d\mid e$
    where $d\mid e$ means $d_{\rho}\mid e_{\rho}$ for any $\rho\in\Sigma(1)$.
In [18, definition 11.9], Salberger defined a Möbius function:
Definition 2.28.
$\mu:(\mathbf{N}^{*})^{\Sigma(1)}\rightarrow\mathbf{Z}$
such that
$\mathbf{1}_{\mathcal{T}(\mathbf{Z})}=\sum\limits_{d\in(\mathbf{N}^{*})^{\Sigma(1)}}\mu(d)\mathbf{1}_{d.\mathbf{Z}^{\Sigma(1)}}.$
For a prime number $p$, we denote by $\sum\limits_{d}^{(p)}$ the sum over the $\Sigma(1)$-tuple of the form $d=(p^{e_{\rho}})_{\rho\in\Sigma(1)}$ where $(e_{\rho})_{\rho\in\Sigma(1)}$ belongs to $\mathbf{N}^{\Sigma(1)}$. We also define $\Pi(d)=\prod\limits_{\rho\in\Sigma(1)}d_{\rho}$ for any $d\in(\mathbf{N}^{*})^{\Sigma(1)}$. Salberger showed (see [18, 11.14]) the following result:
Proposition 2.29.
For any prime number $p$,
$\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p}))=\sum\limits_{d}^{(p)}\frac{\mu(d)}{\Pi(d)}$
Recall that we denote by $f$ be the smallest positive integer such that there exists $f$ rays of $\Sigma$ not contained in a cone of $\Sigma$. We have the following result from [18, lemma 11.15]:
Proposition 2.30.
The Eulerian product over every prime numbers:
$\prod\limits_{p}\sum\limits_{d}^{(p)}\frac{|\mu(d)|}{\Pi(d)^{s}}$
is absolutely convergent for $\Re(s)>\frac{1}{f}$. In the same way,
$\sum\limits_{d\in(\mathbf{N}^{*})^{\Sigma(1)}}\frac{|\mu(d)|}{\Pi(d)^{s}}$
is convergent and is equal to
$\prod\limits_{p}\sum\limits_{d}^{(p)}\frac{|\mu(d)|}{\Pi(d)^{s}}$
for $\Re(s)>\frac{1}{f}$.
Using this proposition, one can deduce (see [18, lemma 11.19]) the following lemma:
Lemma 2.31.
Let $\epsilon\in]0,1-\frac{1}{f}[$. We have the following upper bounds:
  1. (1)
    $\sum\limits_{\Pi(d)\leqslant Y}|\mu(d)|=O\left(Y^{\frac{1}{f}+\varepsilon}\right)$
  2. (2)
    $\sum\limits_{\Pi(d)\geqslant Y}\frac{|\mu(d)|}{\Pi(d)}=O\left(\frac{1}{Y^{1-\frac{1}{f}-\epsilon}}\right).$
3. Lift of the number of rational points of the toric variety to its universal torsor
The method of lifting the number of points to universal torsors have been used in a particular case in [13], then extended to toric varieties in [18, section 11] and generalised in [15],[17].
We fix $u\in(\text{C}_{\text{eff}}(X)^{\vee})^{\circ}$ and $\mathrm{D}\subset\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ a finite union of compact polyhedrons. Let $B>1$ and write $\mathrm{D}_{B}=\mathrm{D}+\log(B)u$. We shall use that the map
$q:\mathcal{T}\rightarrow X$
is surjective. Using the lift of rational points to the universal torsor, we have by corollary 2.16:
$\sharp X(\mathbf{Q})_{h\in\mathrm{D}_{B}}$ $=\frac{1}{2^{t}}\sharp\{y\in\mathcal{T}(\mathbf{Z})\mid h_{\mathcal{T},\infty}(y)\in\mathrm{D}_{B}\}$
$=\frac{1}{2^{t}}\sharp\left(\mathcal{T}(\mathbf{Z})\right)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$
Using the Möbius inversion, we shall see that estimating $\sharp\left(\mathcal{T}(\mathbf{Z})\right)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$ reduces to estimating:
$\sharp\Bigl{(}\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$
3.1. Estimation of the error term between the number of points of a lattice in a bounded domain and the volume of the domain
In what follows, $\operatorname{Vol}$ means the usual euclidean volume given by the Lebesgue measures. The aim of this section is to relate $\sharp\Bigl{(}\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$ and the volume of $\mathcal{D}_{B}=\{y\in\mathcal{T}(\mathbf{R})\mid h_{\mathcal{T},\infty}(y)\in\mathrm{D}_{B}\}$. To estimate the difference between this two terms, we shall use a result of geometry of numbers due to Davenport [8]. The statement we shall use is from [12, proposition 24]:
Proposition 3.1.
Let $\mathcal{D}$ be a bounded domain, semi-algebraic of $\mathbf{R}^{m}$ which is defined by at most $k$ polynomial inequalities each having degree at most $l$.
Then the number of integral lattice points of $\mathbf{Z}^{m}$ contained in $\mathcal{D}$ is :
$\operatorname{Vol}(\mathcal{D})+O\bigl{(}\max(1,\operatorname{Vol}(\operatorname{proj}_{I}(\mathcal{D}))|I\in\mathfrak{P}(\{1,..,m\}),\sharp 1\leqslant I<m)\bigr{)}$
where $\operatorname{Vol}(\operatorname{proj}_{I}(\mathcal{D}))$ means the euclidean volume of the projection of $\mathcal{D}$ via
$\operatorname{proj}_{I}:$ $\mathbf{R}^{m}\rightarrow\mathbf{R}^{I}$
$(x_{1},..,$ $x_{m})\longmapsto(x_{i})_{i\in I}.$
The constant in the error term depends only on $m$, $k$, and $l$.
Each time we shall use this proposition in this article, the constant $m$, $k$ and $l$ will not vary. We shall apply this theorem to the semi-algebraic bounded set $\Phi(\mathcal{D}_{1})$ where:
  1. (1)
    $\mathcal{D}_{1}=\{y\in\mathcal{T}(\mathbf{R})\mid h_{\mathcal{T},\infty}(y)\in\mathrm{D}\}$
  2. (2)
    $\Phi$ is the multiplication by an element $(a_{\rho})_{\rho\in\Sigma(1)}\in(\mathbf{R}_{+}^{\times})^{\Sigma(1)}$. In other words:
    $\Phi:$ $\mathbf{R}^{\Sigma(1)}\rightarrow\mathbf{R}^{\Sigma(1)}$
    $(z_{\rho})_{\rho\in\Sigma(1)}\longmapsto\left(a_{\rho}.z_{\rho}\right)_{\rho\in\Sigma(1)}$
Hence we shall have $m=\sharp\Sigma(1)$, $k$ depends only on the choice of $\mathrm{D}$, the finite union of compact polyhedrons and $l$ depends on the choice of $\mathrm{D}$ and on the maximum of the degree of the monomial $X^{D}$ for $D\in\mathcal{M}_{k}$ and $k\in\{1,..,t\}$ (see notations from 2.2). By this remark, applying the previous proposition gives that there exists $C_{1}>0$ and $C_{2}>0$, which does not depend on the choice of $\Phi$, such that:
$\left|\sharp(\mathbf{Z}^{\Sigma(1)}\cap\Phi(\mathcal{D}_{1}))-\left(\prod\limits_{\rho\in\Sigma(1)}a_{\rho}\right).\operatorname{Vol}(\mathcal{D}_{1})\right|$
$\leqslant C_{1}+C_{2}.\max\limits_{I\subsetneq\Sigma(1)\mid I\neq\varnothing}\left(\left(\prod\limits_{\rho\in I}a_{\rho}\right).\operatorname{Vol}(\operatorname{proj}_{I}(\mathcal{D}_{1}))\right)$
For the rest of the article, we shall suppose that $B\geqslant 1$.
3.1.1. A compacity result
The key point to estimate the error term is the following statement:
Theorem 3.2.
There exists a constant $C>0$ (depending on the compact subset $\mathrm{D}$) such that for any $y\in\mathcal{D}_{B}$, for any $\rho\in\Sigma(1)$, we have:
$|y_{\rho}|\leqslant CB^{\langle[D_{\rho}],u\rangle}.$
From now on, we shall write by $B^{u}$ the unique element of $T_{\operatorname{NS}}(\mathbf{R})$ given by $\exp(\log(B).u)$. Using lemma 2.7, we get that
$\mathcal{D}_{B}=B^{u}.\mathcal{D}_{1}.$
Then to prove the previous theorem, we have to recall that the action of $B^{u}$ on the coordinate associated to $\rho\in\Sigma(1)$ is given by multiplication by $B^{\langle[D_{\rho}],u\rangle}$ by proposition 2.17 and we can conclude using the following lemma:
Lemma 3.3.
There exists a constant $C>0$ (depending on the compact subset $\mathrm{D}$) such that for any $y\in\mathcal{D}_{1}$, for any $\rho\in\Sigma(1)$, we have:
$|y_{\rho}|\leq C.$
Proof.
We can define a surjective continuous map:
$h_{\mathcal{T},\infty}^{-1}(\{0\})$$\times\mathrm{D}\rightarrow\mathcal{D}$
$(y$ $,v)\longmapsto e^{v}.y$
By corollary 2.19, $h_{\mathcal{T},\infty}^{-1}(\{0\})$ is compact. Hence $\mathcal{D}_{1}$ is the image of a compact by a continuous map so it is compact. There exists a constant $C>0$ such that $\mathcal{D}_{1}\subset B_{|-|_{\infty}}(0,C)$. ∎
We get the following corollary of our theorem:
Corollary 3.4.
For any $I$ non empty subset of $\Sigma(1)$, there exists a constant $C>0$ such that:
$\operatorname{Vol}(\operatorname{proj}_{I}(\mathcal{D}_{B}))\leq CB^{\langle\sum\limits_{\rho\in I}[D_{\rho}],u\rangle}.$
We finish this paragraph by recalling the value of the volume $\operatorname{Vol}(\mathcal{D}_{B})$ and by estimating the cardinal of $\sharp\Bigl{(}\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$.
Theorem 3.5.
$\operatorname{Vol}(\mathcal{D}_{B})=\operatorname{Vol}(\mathcal{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}=2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}.$
Proof.
By remark 2.24, the Lebesgue measure over $\mathbf{R}^{\Sigma(1)}$ restricted to $\mathcal{T}(\mathbf{R})$ is $\omega_{\mathcal{T},\infty}$. So the theorem is a consequence of proposition 2.27. ∎
Recall that as we take $u\in(\text{C}_{\text{eff}}(X)^{\vee})^{\circ}$ i.e for any $L\in\text{C}_{\text{eff}}(X)$, $\langle L,u\rangle>0$, then we have that:
$\max\limits_{I\subsetneq\Sigma(1)\mid I\neq\varnothing}\operatorname{Vol}(\operatorname{proj}_{I}(\mathcal{D}_{B}))=O\left(\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}\right)$
which is negligible compared to $\operatorname{Vol}(\mathcal{D}_{B})$ when $B$ grows. We can then deduce the following theorem using proposition 3.1:
Theorem 3.6.
There exists a constant $C>0$ such that:
$\Big{|}\sharp\Bigl{(}\mathbf{Z}^{\Sigma(1)}\cap T_{\Sigma(1)}(\mathbf{R})\Bigr{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\Big{|}\leqslant C\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}$
3.1.2. Upper bound for the number of points in the complementary of the torus
Applying lemma 3.3, we have that there exists $N>1$ such that $\mathcal{D}_{1}\subset\prod\limits_{\rho\in\Sigma(1)}[-N,N]$. Hence $\mathcal{D}_{B}\subset\prod\limits_{\rho\in\Sigma(1)}\left[-N.B^{\langle[D_{\rho}],u\rangle},N.B^{\langle[D_{\rho}],u\rangle}\right]$, so we have for any $\rho\in\Sigma(1)$,
$\sharp\left(\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\cap\{y_{\rho}=0\}\right)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$ $\leqslant\prod\limits_{\rho^{\prime}\in\Sigma(1)-\{\rho\}}\left(2.\lfloor N.B^{\langle[D_{\rho}],u\rangle}\rfloor+1\right)$
$\leqslant\prod\limits_{\rho^{\prime}\in\Sigma(1)-\{\rho\}}\left(2N.B^{\langle[D_{\rho}],u\rangle}+1\right)$
Hence there exists a constant $C>0$ such that:
$\sharp\left(\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\cap\{y_{\rho}=0\}\right)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}\leqslant C.B^{\langle\sum\limits_{\rho\neq\rho^{\prime}}[D_{\rho^{\prime}}],u\rangle}.$
We get the following theorem:
Theorem 3.7.
$\sharp\left(\mathbf{Z}^{\Sigma(1)}\cap(\mathcal{T}(\mathbf{R})\setminus T_{\Sigma(1)}(\mathbf{R}))\right)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}=O\left(\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}\right)$
Applying the previous theorem and theorem 3.6, we get that:
Corollary 3.8.
There exists a constant $C>0$ such that:
$\Big{|}\sharp\Bigl{(}\mathbf{Z}^{\Sigma(1)}\cap\mathcal{T}(\mathbf{R})\Bigr{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\Big{|}\leqslant C\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
As $\sharp\left(X(\mathbf{Q})\setminus T(\mathbf{Q}))\right)_{h\in\mathrm{D}_{B}}\leqslant\sharp\left(\mathbf{Z}^{\Sigma(1)}\cap(\mathcal{T}(\mathbf{R})\setminus T_{\Sigma(1)}(\mathbf{R}))\right)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$,we can also deduce the following corollary:
Corollary 3.9.
$\sharp\left(X(\mathbf{Q})\setminus T(\mathbf{Q}))\right)_{h\in\mathrm{D}_{B}}=O\left(\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}\right).$
3.2. Extension of our estimation in order to apply the Möbius inversion
Let $d\in(\mathbf{N}^{*})^{\Sigma(1)}$ and write $\Lambda_{d}=\bigoplus\limits_{\rho\in\Sigma(1)}d_{\rho}.\mathbf{Z}$. We shall denote $N_{d}(B)$ the cardinal of the set $\Big{(}\Lambda_{d}\cap T_{\Sigma(1)}(\mathbf{R})\Big{)}_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$. We want to extend the estimation made in theorem 3.6 to $N_{d}(B)$.
Denote by $\varphi:\mathbf{R}^{\Sigma(1)}\rightarrow\mathbf{R}^{\Sigma(1)}$ the diagonal endomorphism given by multiplication by $d_{\rho}$ the coordinate associated to the edge $\rho$. Recall that we have $\operatorname{covol}(\Lambda_{d})=\Pi(d)$. We can interpret $N_{d}(B)$ as follows:
$N_{d}(B)$ =$=\sharp\left\{y\in\mathbf{Z}^{\Sigma(1)}\mid h_{\mathcal{T},\infty}\left(\varphi(y)\right)\in\mathrm{D}_{B}\text{ and }\varphi(y)\in T_{\Sigma(1)}(\mathbf{R})\right\}$
=$=\sharp\left\{y\in\mathbf{Z}^{\Sigma(1)}\mid h_{\mathcal{T},\infty}\left(B^{-u}.\varphi(y)\right)\in\mathrm{D}\text{ and }B^{-u}.\varphi(y)\in T_{\Sigma(1)}(\mathbf{R})\right\}$
We write $\mathcal{D}_{1}=\{y\in T_{\Sigma(1)}(\mathbf{R})\mid h_{\mathcal{T},\infty}(y)\in\mathrm{D}\}$. We can then deduce that
$N_{d}(B)=\sharp\left(\mathbf{Z}^{\Sigma(1)}\cap\Phi(\mathcal{D}_{1})\right)$
where $\Phi(z)=\varphi^{-1}(B^{u}.z)$. Using this description, we begin by giving an upper bound of $N_{d}(B)$:
Proposition 3.10.
For any $B>1$, we have:
$N_{d}(B)=O\left(\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{\Pi(d)}\right).$
Proof.
Again by lemma 3.3, there exists $N>1$ such that $\mathcal{D}_{1}\subset\prod\limits_{\rho\in\Sigma(1)}\left([-N,N]-\{0\}\right)$. Hence $\Phi(\mathcal{D}_{1})\subset\prod\limits_{\rho\in\Sigma(1)}\left(\left[-N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}},N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}\right]-\{0\}\right)$. So we have that:
$N_{d}(B)=\sharp(\mathbf{Z}^{\Sigma(1)}\cap\Phi(\mathcal{D}_{1}))\leqslant 2^{\Sigma(1)}.\prod\limits_{\rho\in\Sigma(1)}\lfloor N.\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}\rfloor\leqslant(2.N)^{\sharp\Sigma(1)}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{\Pi(d)}.$
Remark 3.11.
One may remark using the previous proof that $N_{d}(B)=0$ when there exists $\rho\in\Sigma(1)$ such that $N.B^{\langle[D_{\rho}],u\rangle}<d_{\rho}$.
Now we shall later need a better estimation of $N_{d}(B)$ when $\Pi(d)\leqslant B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}$. It is given by the following theorem:
Theorem 3.12.
There exists a constant $C>0$ (depending on the choice of $\mathrm{D})$ such that:
$\left|N_{d}(B)-\frac{1}{\Pi(d)}.2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
Proof.
Denote by $\Phi_{I}:\mathbf{R}^{I}\rightarrow\mathbf{R}^{I}$ the morphism which maps
$(x_{\rho})_{\rho\in I}\mapsto(\frac{B^{\langle[D_{\rho}],u\rangle}}{d_{\rho}}.x_{\rho})_{\rho\in I}.$
Then we may remark that:
$\operatorname{proj}_{I}\circ\Phi=\Phi_{I}\circ\operatorname{proj}_{I}.$
Hence we get, using this, that there exists $C>0$ (which depends only on the compact subset $\mathrm{D}$) such that for any non empty set $I\subsetneq\Sigma(1)$:
$\operatorname{Vol}(\operatorname{proj}_{I}(\Phi(\mathcal{D}))\leqslant C.\frac{B^{\langle\sum\limits_{\rho\in I}[D_{\rho}],u\rangle}}{\prod\limits_{\rho\in I}d_{\rho}}\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
Applying proposition 3.1 finishes the proof of the lemma.
3.3. The Möbius inversion
We recall that we take $\mathrm{D}$ a finite union of compact polyhedrons of $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$, $u\in(\text{C}_{\text{eff}}(X)^{\vee})^{\circ}$ and $B>1$.
Using the definition of 2.28, we have the following proposition:
Proposition 3.13.
$\sharp(\mathcal{T}(\mathbf{Z})\cap T_{\Sigma(1)}(\mathbf{R}))_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}=\sum\limits_{d\in(\mathbf{N}^{*})^{\Sigma(1)}}\mu(d).N_{d}(B)$
Proof.
Firstly, by remark 3.11, the sum is finite. Now we have:
$\sharp(\mathcal{T}(\mathbf{Z})\cap T_{\Sigma(1)}(\mathbf{R}))_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$ $=2^{\sharp\Sigma(1)}.\sum\limits_{y\in(\mathbf{N}^{*})^{\Sigma(1)}}\mathbf{1}_{h_{\mathcal{T},\infty}^{-1}(\mathrm{D}_{B})}(y).\mathbf{1}_{\mathcal{T}(\mathbf{Z})}(y)$
$=2^{\sharp\Sigma(1)}.\sum\limits_{y\in(\mathbf{N}^{*})^{\Sigma(1)}}\mathbf{1}_{h_{\mathcal{T},\infty}^{-1}(\mathrm{D}_{B})}(y).\sum\limits_{d\in(\mathbf{N}^{*})^{\Sigma(1)}}\mu(d).\mathbf{1}_{d.\mathbf{Z}^{\Sigma(1)}}(y)$
$=\sum\limits_{d\in(\mathbf{N}^{*})^{\Sigma(1)}}\mu(d).\left(2^{\sharp\Sigma(1)}.\sum\limits_{y\in(\mathbf{N}^{*})^{\Sigma(1)}}\mathbf{1}_{h_{\mathcal{T},\infty}^{-1}(\mathrm{D}_{B})}(y).\mathbf{1}_{d.\mathbf{Z}^{\Sigma(1)}}(y)\right)$
$=\sum\limits_{d\in(\mathbf{N}^{*})^{\Sigma(1)}}\mu(d).N_{d}(B).$
We shall show the following theorem:
Theorem 3.14.
Let $\varepsilon\in]0,1-\frac{1}{f}[$. There exists a constant $C>0$ such that:
$\left|\sharp(\mathcal{T}(\mathbf{Z})\cap T_{\Sigma(1)}(\mathbf{R}))_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p})).\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|$
$\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{f}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
Proof.
Recall that by propositions 2.29 and 2.30, we have that:
$\sum\limits_{d}\frac{\mu(d)}{\Pi(d)}=\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p})).$
Hence using the previous proposition, we get:
$\left|\sharp(\mathcal{T}(\mathbf{Z})\cap T_{\Sigma(1)}(\mathbf{R}))_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p})).\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|$
$=$ $\left|\sum\limits_{d}\mu(d)\left(N_{d}(B)-2^{t}.\frac{1}{\Pi(d)}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right)\right|$
$\leqslant$ $\sum\limits_{\Pi(d)\leqslant B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}|\mu(d)|.\left|N_{d}(B)-2^{t}.\frac{1}{\Pi(d)}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|$
$+$ $\sum\limits_{\Pi(d)\geqslant B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}|\mu(d)|.\left|N_{d}(B)-2^{t}.\frac{1}{\Pi(d)}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|.$
When $\Pi(d)\leqslant B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}$ we use corollary 3.12, we have that there exists a constant $C>0$ such that:
$\left|N_{d}(B)-\frac{1}{\Pi(d)}.2^{t}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
Let $\varepsilon\in]0,1-\frac{1}{f}[$. Then by lemma 2.31, we get that:
$\sum\limits_{\Pi(d)\leqslant B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}|\mu(d)|.\left|N_{d}(B)-2^{t}.\frac{1}{\Pi(d)}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|$
$\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.B^{(\frac{1}{f}+\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}.$
Now by proposition 3.10, there exists $C^{\prime}>0$ such that:
$N_{d}(B)\leqslant C^{\prime}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{\Pi(d)}.$
Hence we get using the second assertion of lemma 2.31 that there exists $C^{\prime\prime}>0$ such that:
$\sum\limits_{\Pi(d)\geqslant B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}|\mu(d)|.\left|N_{d}(B)-2^{t}.\frac{1}{\Pi(d)}.\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|$
$\leqslant C^{\prime\prime}.B^{\langle\omega_{X}^{-1},u\rangle}.\sum\limits_{\Pi(d)\geqslant B^{\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}\frac{|\mu(d)|}{\Pi(d)}$
$\leqslant C^{\prime\prime}.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{f}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
This concludes our proof. ∎
Applying theorem 3.7, we get:
Corollary 3.15.
Let $\varepsilon\in]0,1-\frac{1}{f}[$. There exists a constant $C>0$ such that:
$\left|(\mathcal{T}(\mathbf{Z}))_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}-2^{t}.\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p})).\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{f}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}$
3.4. Conclusion
Finally we can establish the theorem 1.14:
Theorem 3.16.
Let $\epsilon\in]0,1-\frac{1}{f}[$. Then we have:
$\sharp(X(\mathbf{Q}))_{h\in\mathrm{D}_{B}}\underset{B\rightarrow+\infty}{=}\nu(\mathrm{D}_{1}).\tau(X).B^{\langle\omega_{X}^{-1},u\rangle}\left(1+O\left(B^{-(1-\frac{1}{f}-\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}\right)\right)$
where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 1.8.
Proof.
Let us recall that $\sharp X(\mathbf{Q})_{h\in\mathrm{D}_{B}}=\frac{1}{2^{t}}\sharp\left(\mathcal{T}(\mathbf{Z})\right)_{h_{\mathcal{T},\infty}\in\mathrm{D}_{B}}$. Applying corollary 3.15, we get that there exists a constant $C>0$ such that:
$\left|(\sharp X(\mathbf{Q})_{h\in\mathrm{D}_{B}}-\left(\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p}))\right).\mu_{X,\infty}(X(\mathbf{R})).\nu(\mathrm{D}_{1})B^{\langle\omega_{X}^{-1},u\rangle}\right|\leqslant C.\frac{B^{\langle\omega_{X}^{-1},u\rangle}}{B^{(1-\frac{1}{f}-\varepsilon)\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}}.$
Moreover by remark 2.21 and theorem 2.25, we get:
$\left(\prod\limits_{p}\omega_{\mathcal{T},p}(\mathcal{T}(\mathbf{Z}_{p}))\right).\mu_{X,\infty}(X(\mathbf{R}))=\tau(X)$
hence we have the result stated. ∎